Duane drove at a steady speed for 4 hours on the turnpike. He then slowd his traveling speed by 12 miles per hour for traveling on secondary roads. The entire trip took 7 hours and covered 398 miles. What was his speed on the turnpike?
Answer by mananth(16946)

let speed on turnpike be x mph
distance traveled = 4x
..
speed on secondary road = x-12
distance traveled = 398-4x
..
time = 398-4x / x-12
..
398-4x / x-12 + 4 = 7
398-4x / x-12 = 3
3(x-12) = 398-4x
3x-36= 398- 4x
7x = 398 +36
7x=434
x= 62 mph on turnpike